# Can Large Language Models Analyze Graphs like Professionals? A Benchmark, Datasets and Models

### v2

@@ 3 Benchmark
Figure 1: The pipeline of ProGraph benchmark construction.
@@ 3.1 Tasks
Category 1 32 240 25 15 257 55
Category 2 7 115 7 25 43 111
Category 3 0 30 0 16 0 46
@@ 4 Datasets and Models
Figure 2: The pipeline of LLM4Graph dataset construction and corresponding model enhancement.
@@ 6 Analysis
In this section, we present a comprehensive analysis of the performance of different models on the ProGraph. By grouping the benchmark based on categories, question types, and answer difficulty, we aim to provide a granular exploration of the strengths and limitations of these models in handling graph analysis tasks. We also present the types of compilation errors made by different models.
@@ 6.1 Performance Analysis on Different Benchmark Groupings
To gain a deeper understanding of the capabilities of LLMs and fine-tuned smaller models presented in Section 5.3, we analyze their performances on ProGraph from three different perspectives: task category, question type and answer difficulty.
@@
(b) Question Type
(c) Answer Difficulty
@@
Question Type. We further examine the model performance based on different question types, i.e., true/false, calculation, drawing, and hybrid. In Figure 5(b), we plot the performance of different models on these question types separately. Mainstream LLMs excel in true/false and drawing types but struggle with calculation and hybrid ones. Fine-tuned smaller models demonstrate improvements across various question types, especially on the complex hybrid type, indicating the effectiveness of our proposed enhancement strategies.
Answer Difficulty. Lastly, we divide the ProGraph into two levels of difficulty: easy (involving only one API) and hard (involving multiple APIs). As shown in Figure 5(c), mainstream closed-source large models perform well on easy-level problems, with accuracies generally approaching or exceeding 50%. However, their performance significantly deteriorates on hard-level ones, with the highest accuracy reaching only 32.5%. This observation suggests that mainstream LLMs have limitations when the number of required APIs increases. In contrast, our fine-tuned models demonstrate significantly higher accuracy on hard-level problems, approaching or exceeding 40%, yielding approximately an improvement of 8% compared to the best closed-source LLM. Note that LLM4Graph only contains data involving one API. Still, the models fine-tuned on LLM4Graph show strong generalizability on problems requiring multiple APIs.
@@ 8 Acknowledgements
This work is supported in part by the National Natural Science Foundation of China (No.62192784) and Young Elite Scientists Sponsorship Program (No.2023QNRC001) by CAST.

### v3

@@ 3 Benchmark
Figure 1: The pipeline of ProGraph benchmark construction. We first collect the API documents of 6 frequently-used graph reasoning Python libraries. Then human annotators read the API information, use random graph generators to generate graph data, write questions based on API information, and give the results and the corresponding python codes, forming the Original QA pairs. Finally, GPT-4 rephrases abstract questions and integrates them into real-world scenarios.
@@ 3.1 Tasks
Basic Graph Theory 32 240 25 15 257 55
Graph Statistical Learning 7 115 7 25 43 111
Graph Embedding 0 30 0 16 0 46
@@ 4 Datasets and Models
Figure 2: The pipeline of LLM4Graph dataset construction and corresponding model enhancement. We build the LLM4Graph dataset to improve the capabilities of LLMs in solving graph reasoning tasks through API calls. For the document dataset, we collect API documentation from the Internet. For code dataset, we automatically generate questions and corresponding codes with GPT-4 and API documentation. And for “doc+code” dataset, inspired by CoT reasoning, we add thought for each question, and combine the API documentation and the code dataset for construction.
@@ 5 Experiment
5.1 Experiment Settings and Baselines
In this work, we fine-tune two open-source models, Llama-3-8B-Instruct [1] and Deepseek-Coder-7B-Instruct-v1.5 [18], using the LLM4Graph dataset. For all models, we use alignment-handbook framework [53], set the learning rate, epochs, and max length as 2e-4, 4, and 4096, running on NVIDIA 8*A100 GPUs (40GB). The training batch size is set to 1 and evaluation batch size is set to 4. For testing, we set temperature as 0 and maximum output tokens as 4096, ensuring the stable and reasonable generation.
As our baselines, we employ models without RAG, including the GPT series (GPT-3.5, GPT-4 turbo, GPT-4o) [36], Claude 3 series (Haiku, Sonnet, Opus) [3], and Gemini Pro (1.0 and 1.5) [42], along with non-fine-tuned open-source smaller models Llama-3-8B-Instruct [1] and Deepseek-Coder-7B-Instruct-v1.5 [18]. The detailed results are presented in Table  5.
@@ 6 Analysis
In this section, we present a comprehensive analysis of the performance of different models on the ProGraph. By grouping the benchmark based on categories, answer difficulties, and question type, we aim to provide a granular exploration of the strengths and limitations of these models in handling graph analysis tasks. We also present the types of compilation errors made by different models.
@@ 6.1 Performance Analysis on Different Benchmark Groupings
To gain a deeper understanding of the capabilities of LLMs and fine-tuned smaller models presented in Section 5.4, we analyze their performances on ProGraph from three different perspectives: task category, answer difficulty and question type.
@@
(b) answer difficulty
(c) question type
@@
Answer Difficulty. We further examine the model performance based on different answer difficulties, i.e., true/false, calculation, drawing, and hybrid. In Figure 5(b), we plot the performance of different models on these answer difficulties separately. Mainstream LLMs excel in true/false and drawing types but struggle with calculation and hybrid ones. Fine-tuned smaller models demonstrate improvements across various answer difficulties, especially on the complex hybrid type, indicating the effectiveness of our proposed enhancement strategies.
Question Type. Lastly, we divide the ProGraph into two levels of difficulty: easy (involving only one API) and hard (involving multiple APIs). As shown in Figure 5(c), mainstream closed-source large models perform well on easy-level problems, with accuracies generally approaching or exceeding 50%. However, their performance significantly deteriorates on hard-level ones, with the highest accuracy reaching only 32.5%. This observation suggests that mainstream LLMs have limitations when the number of required APIs increases. In contrast, our fine-tuned models demonstrate significantly higher accuracy on hard-level problems, approaching or exceeding 40%, yielding approximately an improvement of 8% compared to the best closed-source LLM. Note that LLM4Graph only contains data involving one API. Still, the models fine-tuned on LLM4Graph show strong generalizability on problems requiring multiple APIs.
@@ 8 Acknowledgements
This work is supported by the National Natural Science Foundation of China (No.62192784, 62236004), the National Key R&D Program of China (No.2022ZD0116312), Young Elite Scientists Sponsorship Program (No.2023QNRC001) by CAST, and Tsinghua University Initiative Scientific Research Program.